# Entropy-Regularized Process Reward Model

### v1

@@ Abstract
Large language models (LLMs) have shown promise in performing complex multi-step reasoning, yet they continue to struggle with mathematical reasoning, often making systematic errors. A promising solution is reinforcement learning (RL) guided by reward models, particularly those focusing on process rewards, which score each intermediate step rather than solely evaluating the final outcome. This approach is more effective at guiding policy models towards correct reasoning trajectories. In this work, we propose an entropy-regularized process reward model (ER-PRM) that integrates KL-regularized Markov Decision Processes (MDP) to balance policy optimization with the need to prevent the policy from shifting too far from its initial distribution. We derive a novel reward construction method based on the theoretical results. Our theoretical analysis shows that we could derive the optimal reward model from the initial policy sampling. Our empirical experiments on the MATH and GSM8K benchmarks demonstrate that ER-PRM consistently outperforms existing process reward models, achieving 1% improvement on GSM8K and 2-3% improvement on MATH under best-of-N evaluation, and more than 1% improvement under RLHF. These results highlight the efficacy of entropy-regularization in enhancing LLMs’ reasoning capabilities. [^1] Our code is available at https://github.com/hanningzhang/ER-PRM.
[…]
Entropy-Regularized Process Reward Model
Hanning Zhang♠∗,   Pengcheng Wang♡∗,   Shizhe Diao,   Yong Lin, Rui Pan, Hanze Dong, Dylan Zhang, Pavlo Molchanov, Tong Zhang University of Illinois Urbana-Champaign University of Toronto NVIDIA Princeton University Salesforce Research {hanning5, ruip4, shizhuo2, tozhang}@illinois.edu pcheng.wang@mail.utoronto.ca, {sdiao, pmolchanov}@nvidia.com yl7690@princeton.edu, hanze.dong@salesforce.com
**footnotetext: Equal Contribution.
@@ 2 Related Work
Mathematical problem solving is crucial for evaluating LLMs’ reasoning abilities. A common approach is Chain-of-Thought (CoT) prompting, which guides LLMs to solve problems step-by-step (Wei et al., 2023; Zhou et al., 2022; Zhu et al., 2022; Tong et al., 2024). However, general chatbots still struggle with mathematical reasoning, as shown by benchmarks (Cobbe et al., 2021; Hendrycks et al., 2021). To improve performance, several works propose generating synthetic data and fine-tuning LLMs, including Meta-MathQA (Yu et al., 2023), MAmmoTH (Yue et al., 2023), MMIQC (Liu and Yao, 2024), ToRA (Gou et al., 2023), Open-MathInstruct (Toshniwal et al., 2024b), and Dart-Math (Tong et al., 2024). These datasets use teacher models to generate and filter responses, ensuring only correct answers are retained, a process known as rejection sampling or best-of-n sampling (Dong et al., 2023; Yuan et al., 2023). All these emphasize stepwise reasoning for improved task performance.
@@
In mathematical reasoning, reward models in RLHF are not trained via pairwise comparisons due to the availability of ground-truth answers (Bai et al., 2022; Ouyang et al., 2022; Dong et al., 2024). These models are generally divided into Outcome Reward Models (ORM) and Process Supervision Reward Models (PRM). ORM evaluates the overall quality of a response, while PRM provides feedback at each step of the reasoning process. The success of PRM approaches, like Lightman et al. (2023), has fueled interest in studying PRM construction. However, PRMs require extensive human annotation, which is challenging for open-source efforts. To mitigate this, Math-Shepherd (Wang et al., 2024) proposed a method for efficiently acquiring process reward labels by generating parallel completions and using trajectory ratios to evaluate correctness. Building on this, studies have incorporated process supervision and MCTS to improve RLHF pipelines, such as stepwise DPO with MCTS (Chen et al., 2024b), an MCTS-based PRM framework (Luo et al., 2024), MCTS Self-Refine (Zhang et al., 2024), and process generation from final answers (Chen et al., 2024a). These efforts have significantly enhanced process supervision efficiency, motivating the use of MCTS for high-quality data generation.
@@ 3.2 Multi-step Reasoning
We refer to multi-step reasoning as over partial chains by adopting these definitions in appendix H, and our partial reward score assesses the correctness for one step beyond the current partial chain context. Let $a^{[l]}=[a^{1},\dots,a^{l}]$, be partial reasoning chain up to step $l$, and $a^{-[l]}=[a^{l+1},\dots,a^{L}]$ be the completion of the partial reasoning chain from step $l+1$. We define the intermediate reward by $e^{\eta r(a^{[l]},x)}=\mathbb{E}_{a^{-[l]}\sim\pi_{0}}e^{\eta r(a,x)}$. Thus, we have
@@
and we obtain
$\pi_{*}(a^{[l]}|x)=\frac{\pi_{0}(a^{[l]}|x)e^{\eta r(a^{[l]},x)}}{\mathbb{E}_{a^{[l]}\sim\pi_{0}}e^{\eta r(a^{[l]},x)}}.$
@@
An important property of this formulation is that our process reward model can be computed by using the reference policy $\pi_{0}$ to generate the completion. In comparison, in traditional RL, the reward depends on the optimal policy that generates the completion. Therefore in the traditional RL, one has to learn reward and policy simultaneously. This is not necessary in our entropy-based approach. As one can see, equation 1 employs soft optimism over paths generated by the reference policy. More generally, the reward can be computed using any policy including the optimal policy. In this case, the equivalent formula becomes
@@
We note that the process reward has two formulations: when completions are generated from the initial policy $\pi_{0}$, it is soft-max, and when completions are generated from the optimal policy, it is soft-min. If one trains LLM and reward model simultaneously using examples generated by the updated LLM, then one needs to gradually turn from soft max to soft min during the training process.
@@ 4 Experimental Settings
Process Reward Model
We use two widely adopted mathematical datasets, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) in our experiments. We generate and label our process reward data using automatic labeling with entropy regularization, similar to Math-Shepherd (Wang et al., 2024), inspired by Monto Carlo Tree Search (MCTS). In the automatic labeling, the quality of a step is defined as the potential to deduce the final correct answer. For each problem in the training dataset, we use the generator to sample 15 solutions with step-by-step rationale for automatic labeling to increase the diversity of data. For each step, we use the completer to sample 16 paths to reach the final answer. By evaluating the number of correct final answers, the quality of the current step could be reflected. The process reward score for each step is calculated by Equation 1, using initial policy $\pi_{0}$, where $y(a,x)=1$ if the step reaches the correct answer otherwise $y(a,x)=0$.
@@
We use Mistral-7b (Jiang et al., 2023) fine-tuned on MetaMath (Yu et al., 2024) and deepseek-math-7b-instruct (Shao et al., 2024) in our experiments. We use them as the generator and the completer to collect data. For each model, we have around 260K data and 1.5 million step annotations. We train our reward models using Llama-3.1-8B (Dubey et al., 2024). The reward models trained on data from Mistral and deepseek are denoted as Mistral-data Models and deepseek-data Models respectively. The hyper-parameter settings are discussed in Appendix B.
@@
We use the whole test set of GSM8K, and the MATH500 introduced in Lightman et al. (2023) for evaluation. We use the best-of-N strategy to evaluate our reward models. For each problem in the test set, we randomly sample N solutions from the Mistral-MetaMath-7b (Yu et al., 2023) and the deepseek-math-7b-instruct (Shao et al., 2024) generators. Each step in each response is given a score by the reward model. We select the response with the highest score as the final answer. The score for each response is represented using the lowest score among all the steps.
@@
To demonstrate the effectiveness of our method, we compare our models with the hard-label process reward model (Hard-label PRM), vanilla soft-label process reward model (Soft-label PRM), and outcome reward model (ORM). For Hard-label PRM, one reasoning step will be labeled good as long as it can reach the final correct answer. For Soft-label PRM, the step score is approximated by the probability of reaching the final correct answer. For the ORM, it will assign an overall score for the entire mathematical reasoning. The reward calculations are shown and compared in Figure 1. We train the Hard-label PRM, Soft-label PRM, and ORM using the next token prediction, also known as auto-regressive training. We train the PRM to predict the probability of a special token ${}^{\prime}+^{\prime}$ after each reasoning step, and ORM to predict after the entire response. The reward models are trained on these tokens with cross-entropy loss between the prediction and the ground-truth reward The prediction scores would be in the range of $(0,1)$. The dataset format is similar to Math-Shepherd (Wang et al., 2024). We also explore regression training where the reward model is trained to predict a real number as the score given the question and current steps with mean-square error (MSE) loss. The results for regression training are in Section 6.2. We also explore reward models with various sizes in Appendix D, OOD performance with stronger policy models as generators in Appendix E and Appendix F.
Improving Policy Model with PRM
@@
We also design experiments to improve our policy model Mistral-MetaMath-7b using RLHF where we adopt the rejection sampling fine-tuning strategy (Dong et al., 2023) on the MATH training set and OpenMathInstruct-2 dataset (Toshniwal et al., 2024b). For each question, we sample 16 responses and rank them with the reward model. We select the responses with the highest reward score and fine-tune our policy models on these data. The fine-tuned policy models are evaluated on GSM8K and MATH using Top-1 accuracy under a zero-shot CoT setting. We compare different policy models guided by our reward model and guided by other baseline reward models.
@@ 5 Experimental Results
Soft-Label 78.7 30.1
Hard-Label 79.0 32.0
ER-PRM 79.5 32.8
@@
Soft-Label 82.8 43.2
Hard-Label 82.7 43.6
ER-PRM 83.1 44.2
@@
5.1 Auto-Regressive Training Results
The results of our model compared with several baseline methods are shown in Figure LABEL:fig:mistral-policy-best-of-n and Figure LABEL:fig:deepseek-instruct-policy-best-of-n, which are trained in an auto-regressive way.
When the policy model is Mistral-MetaMath-7b, as N increases, the best-of-N accuracy for most of the models keeps increasing except the Outcome Reward Model (ORM). In particular, ER-PRM outperforms other baselines in all evaluation settings. Our model demonstrates consistent improvements, with more substantial gains observed on the more challenging datasets. Concretely, on MATH, which presents harder questions compared with GSM8K, the gap widens significantly, highlighting our method’s more pronounced effectiveness in tackling complex datasets.
When the models are trained on deepseek data and evaluate on deepseek-math-7b-instruct, the results are similar, where the accuracy is boosted by a large margin. And ER-PRM outperforms Soft-label PRM and Hard-label PRM on GSM8K and MATH500. However, we notice that on MATH500, the increment from $N=4$ to $N=64$ is not very significant. And there are reward hacking problems with large $N$. Similar Observations are also reflected in Section F. Interestingly, we observe a very strong performance of ORM on GSM8K, demonstrating the potential of training a powerful ORM with high-quality data.
We notice that the Hard-label PRM also demonstrates similar strong performance for best-of-N evaluation, and is effective as Soft-label PRM in some cases, especially when N is large. We attribute this to the semi-supervised learning process during the hard-label training. The labels during the Markov Decision Process are only an estimation of the process reward instead of a precise value. The Hard-label setting is a way to reduce the noise during the training.
We note that here the Hard-label setting adopts the same labeling strategy with Math-Shepherd (Wang et al., 2024), trained with the same base model as ER-PRM to ensure a fair comparison. Our model outperforms Math-Shepherd in all settings on GSM8K and MATH, demonstrating the ability to better rank and select candidate responses. It shows that ours are intrinsically competent reward models compared to the open-source models with high consistency and robustness.
5.2 RLHF Results
We choose to verify the effectiveness of our auto-regressive Entropy-Regularized Process Reward Model (ER-PRM) by comparing it against the Soft-Label PRM and Hard-Label PRM in RLHF via RAFT algorithm (Dong et al., 2023). We use the Mistral-MetaMath-7b and the deepseek-Math-7B-Instruct as the initial baseline policy models, and the reward models trained on Mistral data and deepseek data.
@@ 5.2 RLHF Results
During initial tests, we observed that the baseline policy model exhibited signs of overfitting to the training data from GSM8K and MATH, resulting in only marginal performance gains after RLHF on the original training set, particularly for deepseek-Math-7B-Instruct. To address this, we incorporated out-of-domain data from OpenMathInstruct-2 (Toshniwal et al., 2024a), augmenting GSM8K and MATH with 12,500 additional new questions each. For each question, the initial policy model generated 16 candidate rationales, which were then evaluated by both ER-PRM and baseline PRMs to select the highest-scoring rationale as training data for RLHF.

### v2

@@ Abstract
Large language models (LLMs) have shown promise in performing complex multi-step reasoning, yet they continue to struggle with mathematical reasoning, often making systematic errors. A promising solution is reinforcement learning (RL) guided by reward models, particularly those focusing on process rewards, which score each intermediate step rather than solely evaluating the final outcome. This approach is more effective at guiding policy models towards correct reasoning trajectories. In this work, we propose an entropy-regularized process reward model (ER-PRM) that integrates KL-regularized Markov Decision Processes (MDP) to balance policy optimization with the need to prevent the policy from shifting too far from its initial distribution. We derive a novel reward construction method based on the theoretical results. Our theoretical analysis shows that we could derive the optimal reward model from the initial policy sampling. Our empirical experiments on the MATH and GSM8K benchmarks demonstrate that ER-PRM consistently outperforms existing process reward models, achieving 1% improvement on GSM8K and 2-3% improvement on MATH under best-of-N evaluation, and more than 1% improvement under RLHF. These results highlight the efficacy of entropy-regularization in enhancing LLMs’ reasoning capabilities.
@@ 2 Related Work
Mathematical problem solving is crucial for evaluating LLMs’ reasoning abilities. A common approach is Chain-of-Thought (CoT) prompting, which guides LLMs to solve problems step-by-step (Wei et al., 2023; Zhou et al., 2022; Zhu et al., 2022; Tong et al., 2024). However, general chatbots still struggle with mathematical reasoning, as shown by benchmarks (Cobbe et al., 2021; Hendrycks et al., 2021). To improve performance, several works propose generating synthetic data and fine-tuning LLMs, including Meta-MathQA (Yu et al., 2023), MAmmoTH (Yue et al., 2023), MMIQC (Liu & Yao, 2024), ToRA (Gou et al., 2023), Open-MathInstruct (Toshniwal et al., 2024b), and Dart-Math (Tong et al., 2024). These datasets use teacher models to generate and filter responses, ensuring only correct answers are retained, a process known as rejection sampling or best-of-n sampling (Dong et al., 2023; Yuan et al., 2023). All these emphasize stepwise reasoning for improved task performance.
@@
In mathematical reasoning, reward models in RLHF are not trained via pairwise comparisons due to the availability of ground-truth answers (Bai et al., 2022; Ouyang et al., 2022; Dong et al., 2024). These models are generally divided into Outcome Reward Models (ORM) and Process Supervision Reward Models (PRM). ORM evaluates the overall quality of a response, while PRM provides feedback at each step of the reasoning process. The success of PRM approaches, like Lightman et al. (2023), has fueled interest in studying PRM construction. However, PRMs require extensive human annotation, which is challenging for open-source efforts. To mitigate this, Math-Shepherd (Wang et al., 2024) proposed a method for efficiently acquiring process reward labels by generating parallel completions and using trajectory ratios to evaluate correctness. This has been proved as effective as human annotation to generate high-quality datasets. Building on this, studies have incorporated process supervision and MCTS to improve RLHF pipelines, such as stepwise DPO with MCTS (Chen et al., 2024b), an MCTS-based PRM framework (Luo et al., 2024), MCTS Self-Refine (Zhang et al., 2024), and process generation from final answers (Chen et al., 2024a). These efforts have significantly enhanced process supervision efficiency, motivating the use of MCTS for high-quality data generation.
@@ 3.2 Multi-step Reasoning
We refer to multi-step reasoning as over partial chains by adopting these definitions in appendix D, and our partial reward score assesses the correctness for one step beyond the current partial chain context. Let $a^{[l]}=[a^{1},\dots,a^{l}]$, be partial reasoning chain up to step $l$, and $a^{-[l]}=[a^{l+1},\dots,a^{L}]$ be the completion of the partial reasoning chain from step $l+1$. We define the intermediate reward by $e^{\eta r(a^{[l]},x)}=\mathbb{E}_{a^{-[l]}\sim\pi_{0}}e^{\eta r(a,x)}$. Thus, we assume we have prompt $x$ and the first $a^{[l]}$ steps, we get the derivation for the remaining $a^{-[l]}$:
@@
and given prompt $x$ and we obtain the derivation for $a^{[l]}$:
$\pi_{*}(a^{[l]}|x)=\frac{\pi_{0}(a^{[l]}|x)e^{\eta r(a^{[l]},x)}}{\mathbb{E}_{a^{[l]}\sim\pi_{0}}e^{\eta r(a^{[l]},x)}}.$ (1)
@@
An important property of this formulation is that our process reward model can be computed by using the reference policy $\pi_{0}$ to generate the completion. In comparison, in traditional RL, the reward depends on the optimal policy that generates the completion. Therefore in the traditional RL, one has to learn reward and policy simultaneously. This is not necessary in our entropy-based approach. As one can see, equation 2 employs soft optimism over paths generated by the reference policy. More generally, the reward can be computed using updated policy including the optimal policy. In this case, the equivalent formula becomes
@@
We note that the process reward has two formulations: when completions are generated from the initial policy $\pi_{0}$, it is soft-max, and when completions are generated from the optimal policy, it is soft-min. If one trains LLM and reward model simultaneously using examples generated by the updated LLM, then one needs to gradually turn from soft max to soft min during the training process. For ease of understanding, we provide intuitive explanations of our approach’s soft-max and soft-min perspectives in Appendix E, along with a detailed analysis of the theoretical advantages of entropy regularization for process rewards.
@@ 4 Experimental Settings
We incorporated a figure in Appendix F to further clarify our experiment settings.
Process Reward Construction
We conduct experiments using two widely adopted mathematical datasets: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). Our process reward data is generated and labeled through an automatic labeling approach with entropy regularization, similar to Math-Shepherd (Wang et al., 2024), and is inspired by Monte Carlo Tree Search (MCTS).
In this automatic labeling process, the quality of each step is defined by its potential to deduce the correct final answer. To enhance data diversity, we use a generator to sample 15 solutions per problem, each including step-by-step reasoning. For each step, we employ a completer to generate 16 possible paths leading to the final answer. The quality of a given step is then assessed based on the number of sampled paths that produce the correct final answer. The process reward score for each step is computed using Equation 2, based on the initial policy $\pi_{0}$. Specifically, the reward function $r(a,x)$ assigns a value of 1 if the step leads to the correct final answer and 0 otherwise.
Process Reward Sampling
@@
We conduct experiments using Mistral-7B (Jiang et al., 2023), fine-tuned on the MetaMath dataset (Yu et al., 2024) which we denote as Mistral-MetaMath-7B, and DeepSeek-math-7B-instruct (Shao et al., 2024). These models serve as the generator and completer for data collection, using the method introduced in the paragraph above. For each model, we collect approximately 260K data points and 1.5 million step annotations. Our reward models are trained using Llama-3.1-8B (Dubey et al., 2024).
The reward models are trained on the following datasets:
  • $\bullet$
    Mistral-data: Process reward training data collected using Mistral-MetaMath-7B (Yu et al., 2024) as both the generator and completer.
  • $\bullet$
    DeepSeek-data: Process reward training data collected using DeepSeek-math-7B-instruct (Shao et al., 2024) as both the generator and completer.
In our main experiments (Section 5), we train the reward models using Llama-3.1-8B on both Mistral-data and DeepSeek-data. Additionally, in Section 6.1, we investigate reward model training at different scales by using Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct (Dubey et al., 2024) on Mistral-data.
Process Reward Model Training
The reward models are trained with a global batch size of 64, a learning rate of 1e-6, a maximum sequence length of 512, and a single epoch using four H100 GPUs. We perform hyperparameter tuning on the learning rate, exploring values from $5e^{-6}$$2e^{-6}$$1e^{-6}$$5e^{-7}$}. During the automatic labeling process, we set $\eta=2$ for Mistral data, and $\eta=10$ for DeepSeek data in our main experiments. We also conduct hyperparameter tuning for $\eta$ across the values {$0.1$, 11, 22, 55, 88, 1010, 1515}.
Process Reward Model Evaluation
@@
We evaluate our reward models using the full test set of GSM8K and the MATH500 dataset introduced by Lightman et al. (2023). Our evaluation follows a best-of-N strategy. For each problem in the test set, we randomly sample N solutions from the generators. Each step within a generated response is scored by the reward model. The final answer is selected as the response with the highest overall score. The score for each response is determined by the lowest-scoring step within it.
Baseline Settings and Training Details
To evaluate the effectiveness of our method, we compare our models against three baselines: the hard-label process reward model (Hard-label PRM), the vanilla soft-label process reward model (Soft-label PRM), and the outcome reward model (ORM).
  • $\bullet$
    Hard-label PRM: A reasoning step is labeled as good if it ultimately leads to the correct final answer.
  • $\bullet$
    Soft-label PRM: The step score is approximated by the probability of reaching the correct final answer.
  • $\bullet$
    ORM: Assigns a single overall score to the entire reasoning process indicating correct or not.
@@
The reward calculation methods for these models are illustrated in Figure 1. We train the Hard-label PRM, Soft-label PRM, and ORM using the next token prediction, also known as auto-regressive training. Specifically, PRM is trained to predict the probability of a special token $^{\prime}+^{\prime}$ after each reasoning step, while ORM predicts this token at the end of the response. The reward models are optimized using cross-entropy loss between the predicted and ground-truth rewards, producing scores in the range of $(0,1)$. The dataset format follows Math-Shepherd (Wang et al., 2024). In addition to the next token prediction, we explore regression-based training in Section 6.4. We also explore reward models with various sizes in Section 6.1, and out-of-distribution (OOD) performance with stronger policy models as generators in Section 6.2. For each experiment settings, we repeat 5 times with different seeds and report the average accuracy along with the standard deviation to ensure the reliability of the results.
Policy Optimization with PRM-Guided Reinforcement Learning
@@
We conduct experiments to enhance our policy models, Mistral-MetaMath-7B and DeepSeek-Math-7B-Instruct, using reinforcement learning from human feedback (RLHF). Specifically, we adopt the rejection sampling fine-tuning strategy (Dong et al., 2023) on the MATH training set and the OpenMathInstruct-2 dataset (Toshniwal et al., 2024b). For each question, we generate 16 responses and rank them using our process reward model (PRM). We then select the highest-scoring responses and fine-tune the policy model on this filtered dataset. The fine-tuned policy models are evaluated on GSM8K and MATH using Top-1 accuracy under a zero-shot chain-of-thought (CoT) setting. Additionally, we compare policy models optimized with our reward model against those guided by baseline reward models. We also report the average accuracy and the standard deviation.
@@ 5 Experimental Results
5.1 Auto-Regressive Training Results
(a) Best-of-N evaluation results on GSM8K and MATH500 datasets with Mistral-MetaMath-7b as the generator. The reward models are trained on Mistral-data (GSM8K and MATH).
(b) Best-of-N evaluation results on GSM8K and MATH500 datasets with DeepSeek-math-7b-instruct as the generator. The reward models are trained on DeepSeek-data.
We train the reward model with Llama-3.1-8B as the base model. The results of our model compared with several baseline methods are shown in Figure 3(a) and Figure 3(b), which are trained in an auto-regressive way. The details are also listed in Table 5 and Table 6 in the Appendix.
Mistral-MetaMath-7b as the Generator, Reward Models trained on Mistral-data
When using Mistral-MetaMath-7B as the policy model, we observe that as N increases, the best-of-N accuracy improves for most models. Notably, ER-PRM consistently outperforms all baseline models across all evaluation settings. In general, process reward models (PRMs) outperform outcome reward models (ORMs), demonstrating the effectiveness of stepwise reward modeling. Our method shows consistent improvements, with particularly significant gains on more challenging datasets. Specifically, on MATH, which contains more complex problems than GSM8K, the performance gap widens substantially. This highlights our approach’s superior ability to handle difficult mathematical reasoning tasks.
DeepSeek-math-7b-instruct as the Generator, Reward Models trained on DeepSeek-data
When trained on DeepSeek-data and evaluated on DeepSeek-math-7B-Instruct, the results show a significant accuracy boost. ER-PRM consistently outperforms Soft-label PRM and Hard-label PRM on GSM8K and MATH500. However, we observe that on MATH500, the performance improvement from N=4 to N=64 is less pronounced. Additionally, we identify reward hacking issues when using large N. Similar trends are also observed in Section 6.2, where OpenMath2-Llama3.1-8B is used as the policy model. Interestingly, ORM achieves exceptionally strong performance on GSM8K, suggesting that a well-trained ORM with high-quality data can be highly effective.
We also find that Hard-label PRM exhibits strong performance in best-of-N evaluation. In most cases, especially for large N, it surpasses Soft-label PRM. We attribute this to the semi-supervised nature of the Hard-label training process. Since the Markov Decision Process (MDP) labels are only approximations of the process reward rather than precise values, the Hard-label setting helps reduce noise during training.
We note that here the Hard-label setting adopts the same labeling strategy with Math-Shepherd (Wang et al., 2024), trained with the same base model as ER-PRM to ensure a fair comparison. Our model outperforms Math-Shepherd in nearly all settings on GSM8K and MATH, demonstrating the ability to better rank and select candidate responses. These results highlight the consistency and robustness of our reward models, making them more competent than existing open-source alternatives.
5.2 RLHF Results
To verify the effectiveness of our auto-regressive Entropy-Regularized Process Reward Model (ER-PRM), we compare it against the Soft-Label PRM and Hard-Label PRM within the context of RLHF using the RAFT algorithm (Dong et al., 2023). We employ Mistral-MetaMath-7b and DeepSeek-math-7B-instruct as the initial baseline policy models and adopt the reward models trained on Mistral data and DeepSeek data, respectively.
5.2.1 Manifestation and Mitigation of Overfitting
@@
Soft-Label 78.5±0.3 30.4±0.8
Hard-Label 79.0±0.5 31.5±0.5
ER-PRM 79.7±0.3 32.5±0.4
@@
Soft-Label 82.9±0.4 42.8±0.4
Hard-Label 82.4±0.3 43.8±0.8
ER-PRM 83.4±0.3 44.7±0.7
@@
To clarify our evaluation methodology, all results reported in this paper were evaluated exclusively on the test sets of GSM8K and MATH, never on training data. During initial experiments, we observed that when using only the original GSM8K and MATH training sets for RLHF, the models showed limited generalization to the test sets, suggesting potential overfitting since these models had already been exposed to these datasets during their original supervised fine-tuning. Our analysis confirmed that Mistral-MetaMath-7B was fine-tuned on MetaMathQA (which incorporates GSM8K and MATH problems), while DeepSeek-Math-7B-Instruct explicitly included these datasets in its training corpus.
@@
To address this overfitting challenge while preserving the consistency of model behaviour, we incorporated 12,500 novel questions from the OpenMathInstruct-2 dataset (Toshniwal et al., 2024a) into our RLHF training data to conduct a blended training. This dataset contains synthetic mathematical problems and rationales that are not present in the original training data of our policy models. For each question, the initial policy model generated 16 candidate rationales (the number of candidates was selected by grid search), which were then evaluated by the ER-PRM and baseline PRMs to select the highest-scoring rationale for RLHF training. The OpenMathInstruct-2 data was used only for training, never for evaluation.
5.3 Performance Evaluation